when 1 mole of barium sulfate was dissolved in water it resulted to a H of 0.52 KJ/mol. How much heat is absorbed when 450g if the compound is dissolved in water has zero effect
1
Expert's answer
2020-05-05T03:40:27-0400

Q= m(BaSO4)/Mr(BaSO4)*H=(450/233)*0.52=1 kJ

Answer: 1kJ of heat is absorbed